OpenStudy (anonymous):

Find each of the following. f(x) = |x + 3| g(x) = -x^2 (g∘f)(1)

OpenStudy (anonymous):

If \(g(u)=-u^2,\) and \(u=f(x)=|x+3|,\) then \(g(f(x))=-|x+3|^2\).

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!